# Quantum walks, quantum cellular automata, quantum field theory, symmetry, interactions

### v1

@@ 4 Momentum Space Picture and the Dirac and Maxwell Equations
where $-\pi/\Delta x<k_{X,Y,Z}\leq\pi/\Delta x$. For mathematical simplicity, we assume that the overall size of the lattice is finite, albeit large, with periodic boundary conditions; this choice makes the momentum discrete, and the momentum states renormalizable. If the cubic lattice is $N\times N\times N$ then the normalization factor is $\mathcal{N}=N^{3/2}$. If necessary we can then go to a limit where $N\rightarrow\infty$. In the finite case, the momentum values $k_{X,Y,Z}$ are integer multiples of $2\pi/N\Delta x$.
@@
where the $K_{X,Y,Z}$ are the operators corresponding to the components of the momentum and the $\{\Lambda_{\mathbf{k}}\}$ are the diagonalized forms of the $\{U_{\mathbf{k}}\}$. Using the choices above for $\Delta P_{X,Y,Z}$, we have
@@
where for fermions these are the usual gamma matrices, and for bosons the gamma matrices are the boson versions defined, and the first term is absent, i.e., $\theta=0$. (As we will see, this corresponds to the massless case.) In this representation we see that the Hilbert space decomposes into $D$-dimensional blocks ($D=4$ for fermions, $D=6$ for bosons) labeled by the momentum $\mathbf{k}$. The energy eigenstates ${|\mathbf{k},\lambda_{\mathbf{k},j}\rangle}$ of the quantum walk can be found by diagonalizing the $D\times D$ matrices $U_{\mathbf{k}}=V_{\mathbf{k}}\Lambda_{\mathbf{k}}V^{\dagger}_{\mathbf{k}}$. Since $U_{\mathbf{k}}$ is unitary, the matrices $\Lambda_{\mathbf{k}}$ have eigenvalues of the form $\lambda_{\mathbf{k},j}=e^{i\phi_{\mathbf{k},j}}$.
For fermions, if we go to the long-wavelength/low-mass limit $|\mathbf{k}\Delta x|\ll 1$ and $|\theta|\ll 1$, we get
@@
In terms of this we can get a Dirac equation, which holds in the long-wavelength/low-mass limit:
@@
We can relate the physical space interval $\Delta x$ and the physical time interval $\Delta t$ in terms of the speed of light $c$: $c\equiv\Delta x/\Delta t$. Then, defining the particle mass $\theta/\Delta t\equiv m_{e}c^{2}$, and using $\Delta P_{i}=\gamma_{0}\gamma_{i}$, we obtain the momentum-space representation of the Dirac equation:
@@
For bosons, we get the analogous equation, except with $\Delta P_{i}=\gamma_{0}^{\rm bose}\gamma_{i}^{\rm bose}$. This gives, in the same small $\mathbf{k}$ limit,
@@
with eigenvalues 00 and ±$\pm ik$, respectively, where $k=|\mathbf{k}|=\sqrt{k_{X}^{2}+k_{Y}^{2}+k_{Z}^{2}}$. These correspond to a longitudinal mode, and right- and left-polarized modes, respectively. The Hamiltonian thus has two zero-energy longitudinal eigenstates, and a positive and negative energy eigenstate for each circular polarization. In the free theory these do not mix, so if we consider positive-energy initial states, the dynamics do not take us out of the positive-energy subspace. That is not generally true in the presence of interactions, which we will discuss later.
When we promote the single-particle theory to allow for the creation and annihilation of photons, we’ll find that the field operators, expressed in the usual manner in terms of the creation and annihilation operators and the above eigenvectors, will satisfy the Maxwell equations in the long wavelength limit. To see that here, for the one-particle sector, we have to consider the photon wave function (see Refs. BialynickiBirula ; Darwin32 ):
@@
and the vector identity in Eq. (25), we get
@@
Substituting $-i\boldsymbol{\nabla}$ for $c\mathbf{k}$ and considering the real and imaginary parts separately, we get the two time-dependent Maxwell equations:
@@
The other two Maxwell equations (in the absence of charge and current) arise from the transversality of the positive energy eigenvectors.
@@ 5.1 Fermions
As was shown in MlodinowBrun21 , it is possible to promote the single-particle fermionic theory to a multi-particle QCA that reduces to the usual Dirac field theory in the long-wavelength limit. To accomplish this, one begins with a maximum number of particles, $N_{\rm max}$, and then let $N_{\rm max}$ get arbitrarily large. In other words, one begins with the space
@@
The subspace $\mathcal{A}_{n}$ is the totally antisymmetric subspace with $n$ particles and $\mathcal{H}_{0}$ is the one-dimensional vacuum space ${\rm span}({|\Omega\rangle})$, where ${|\Omega\rangle}={{|\omega\rangle}}^{\otimes N_{\rm max}}$. We can define basis vectors for the subspaces $\mathcal{A}_{n}$ in terms of the “energy eigenstates” of the quantum walk:
@@
The $n$-particle antisymmetrized subspace is thus preserved by the time evolution, and hence so is the full physical subspace $\mathcal{H}_{\rm phys}$: an antisymmetric initial state will evolve to be antisymmetric at all later times. Given this basis, we can define a set of creation and annihilation operators that transform the energy basis vectors into each other:
@@
and that obey the usual anticommutation relations. To avoid redundant counting of the basis vectors, we should establish a standard ordering for the states $\mathbf{k}_{i},j_{i}$. The choice of this ordering is unimportant, so long as it is unambiguous.
@@
A very important observation should be made here. These creation and annihilation operators $a^{\dagger}_{\mathbf{k},j}$ and $a_{\mathbf{k},j}$ are clearly nonlocal operators; so the unitary operator on the right-hand side of Eq. (37) is nonlocal. But the unitary operator $U_{\rm total}$ is itself purely local. It is a collection of local quantum walk unitaries, and can easily be embedded in a quantum cellular automaton. These two sides are equal only on states in the physical subspace. Indeed, the right-hand side has not been defined for states outside that subspace.
@@
using the matrices $V^{\dagger}_{\mathbf{k}}$ from Eq. (16). These vectors have an effective time evolution
@@
where $U_{\mathbf{k}}$ is given by Eq. (16).
What about the maximum number of particles $N_{\rm max}$? This may seem like an unphysical feature of the theory. In the free theory, particle number is conserved, so a solution will generally be confined to a single antisymmetric subspace $\mathcal{A}_{n}$. However, in the interacting case this conservation need not hold.
One way to handle this is to start with a vacuum corresponding to a Dirac sea, in which all the negative-energy states are occupied while all positive-energy states are vacant. Interactions do not create or destroy particles, but may move a particle from a negative- to a positive-energy state, leaving a hole (corresponding to particle-antiparticle pair creation), or move a positive-energy state to a negative-energy hole (corresponding to particle-antiparticle annihilation). In such a construction we would have $N_{\rm max}=2N^{3}$, where $N$ is the lattice size. This means that the Hilbert spaces of the local subsystems at each lattice site are enormously high-dimensional (albeit finite), but the states of the theory remain restricted to the much lower-dimensional physical subspace. In the full space the dynamics of the free theory are fully local; the effective dynamics in the physical subspace obey Fermi statistics.
@@ 5.2 Bosons
We can make a similar construction for free bosons, with $N_{\rm max}$ copies of a bosonic QW, representing distinguishable particles, and then confining ourselves to the completely symmetric subspace. So long as the total number of bosonic particles $N$ is less that $N_{\rm max}$, this system will behave like a bosonic field theory with $N$ excitations, and we can introduce creation and annihilation operators and rewrite the evolution unitary in terms of them, just as we did in the fermionic case.
@@
where $j=X,Y,Z$. As discussed in Section 4 we can recover a Maxwell field if we use a doubled space, where each of the projectors is two-dimensional, for an internal space dimension of 6. To simplify the derivation of the corresponding QCA, we will first consider a non-doubled space, with internal dimension 3, so we can write the projectors as outer products of the eigenstates of the $J_{j}$ matrix:
@@
In other words, the shifts are done in two steps: the $+$ and -- states at neighboring sites are interchanged, and then at each site $+$ and -- are switched back. (The 0 states remain unchanged throughout.)
The advantage of breaking this down into two steps is that each step is a purely local operation that only affects two neighboring sites. This makes it very easy to turn into a QCA that is definitely local. In constructing the QCA, each basis state ${|\mathbf{x},\varepsilon\rangle}$ of the QW corresponds to a harmonic oscillator mode with creation and annihilation operators $a^{\dagger}_{\mathbf{x},\varepsilon},a_{\mathbf{x},\varepsilon}$. Then the evolution unitary becomes
@@
We can see that $\hat{U}_{j}$ is a product of tensor products of unitaries, which act on the local subsystems $\mathbf{x}$ or on non-overlapping sets of modes at neighboring subsets $\mathbf{x}$ and $\mathbf{x}+\Delta x\mathbf{\hat{j}}$. So $\hat{U}_{j}$ is local, and defines a QCA. Since operators on different modes commute, we can rewrite each tensor product of exponentials as the exponential of a sum of terms (where now we understand the creation and annihilation operators as acting as the identity on other modes). Further, we see that the defined QCA preserves particle number (or the number of excitations); and on the single-particle sector, it exactly reproduces a quantum walk.
@@
Each of the operators $\hat{U}_{X,Y,Z}$ has the same form as in Eq. (42). In the quantum walk, we get nontrivial evolution because the projectors $P^{\pm}_{X,Y,Z}$ for the shifts along each axis project onto subspaces that are non-orthogonal to those of the other axes. For the QCA, this means that the creation and annihilation operators in Eq. (42) for one axis $j$ are linear combinations of those for a different axis $j^{\prime}$.
Since all the expressions inside the exponentials are quadratic in creation and annihilation operators, we can rewrite Eq. (42) in terms of a standard set of modes with operators $a^{\dagger}_{\mathbf{x},\ell}$ and $a_{\mathbf{x},\ell}$:
@@
It is easy to check that creation and annihilation operators for different values of $\mathbf{k}$ commute with each other. If we substitute this definition into Eq. (44) the unitary decomposes into a product of factors acting on different momentum modes:
@@
This is most readily done by diagonalizing $A$ and $B$, exponentiating them, and then multiplying them. One can find $C$ by diagonalizing the resulting matrix and taking its log. In the case of the unitary factor $U_{j}$, the matrices $A_{j}$ and $B_{j}$ have elements
@@
These matrices are $3\times 3$, so in general it is not difficult to diagonalize them. For a bosonic QW, we suggested in Sec. 3 that the basis vectors should be eigenvectors of the three spin-1 $J$ matrices given in Eq. (12). This choice for the bosonic QCA would lead to three pairs of $A$ and $B$ matrices:
@@
where we have used the compact notation $c_{x,y,z}=\cos(k_{x,y,z}\Delta x)$ and $s_{x,y,z}=\sin(k_{x,y,z}\Delta x)$. This matrix can, in principle, be diagonalized in closed form, but the expression is dauntingly complicated. Its eigenvalues can also be calculated, though they are also somewhat complicated:
@@
Simple observation shows that the $+$ and -- eigenvalues have opposite phases: $\phi_{-}=-\phi_{+}$. These can be thought of as negative and positive energy modes. We can recover more familiar expressions by going to the long-wavelength limit, where $|\mathbf{k}\Delta x|\ll 1$. In this limit
@@
As discussed in Section 4, we can recover the Maxwell field theory by doubling the internal space. In place of the spin matrices $J_{j}$ we use the $6\times 6$ matrices $\sigma_{Z}\otimes J_{j}$. The derivation above is essentially doubled, giving two zero-energy, two negative-energy and two positive-energy modes for each allowed momentum $\mathbf{k}$. If we confine ourselves to the positive-energy subspace, these two modes correspond to the two allowed helicity states of the electromagnetic field. The two negative-energy modes are like retarded versions of these two modes. The zero-energy modes are static and unaffected by the other modes.
We have seen that the free theories of both fermionic and bosonic QCAs include matching positive and negative energy modes. (And zero-energy modes as well, in the bosonic case.) This turns out to be a general property of QCAs based on their overall symmetry, as discussed in the next section.

### v2

@@ 4 Momentum-Space Picture and the Dirac and Maxwell Equations
where $-\pi/\Delta x<k_{X,Y,Z}\leq\pi/\Delta x$. For mathematical simplicity, we assume that the overall size of the lattice is finite, albeit large, with periodic boundary conditions; this choice makes the momentum discrete and the momentum states renormalizable. If the cubic lattice is $N\times N\times N$, then the normalization factor is $\mathcal{N}=N^{3/2}$. If necessary, we can then go to a limit where $N\rightarrow\infty$. In the finite case, the momentum values $k_{X,Y,Z}$ are integer multiples of $2\pi/N\Delta x$.
@@
where $K_{X,Y,Z}$ are the operators corresponding to the components of the momentum, and $\{\Lambda_{\mathbf{k}}\}$ are the diagonalized forms of $\{U_{\mathbf{k}}\}$. Using the choices above for $\Delta P_{X,Y,Z}$, we have
@@
where, for fermions, these are the usual gamma matrices, and for bosons, the gamma matrices are the boson versions defined, and the first term is absent, i.e., $\theta=0$. (As we will see, this corresponds to the massless case.) In this representation, we see that the Hilbert space decomposes into $D$-dimensional blocks ($D=4$ for fermions, $D=6$ for bosons) labeled by the momentum $\mathbf{k}$. The energy eigenstates ${|\mathbf{k},\lambda_{\mathbf{k},j}\rangle}$ of the quantum walk can be found by diagonalizing the $D\times D$ matrices $U_{\mathbf{k}}=V_{\mathbf{k}}\Lambda_{\mathbf{k}}V^{\dagger}_{\mathbf{k}}$. Since $U_{\mathbf{k}}$ is unitary, the matrices $\Lambda_{\mathbf{k}}$ have eigenvalues of the form $\lambda_{\mathbf{k},j}=e^{i\phi_{\mathbf{k},j}}$.
For fermions, if we go to the long-wavelength/low-mass limit $|\mathbf{k}\Delta x|\ll 1$ and $|\theta|\ll 1$, we obtain
@@
In terms of this, we can obtain a Dirac equation, which holds in the long-wavelength/low-mass limit:
@@
We can relate the physical space interval $\Delta x$ and the physical time interval $\Delta t$ in terms of the speed of light $c$: $c\equiv\Delta x/\Delta t$. Then, defining the particle mass $\theta/\Delta t\equiv m_{e}c^{2}$ and using $\Delta P_{i}=\gamma_{0}\gamma_{i}$, we obtain the momentum-space representation of the Dirac equation:
@@
For bosons, we obtain the analogous equation, except with $\Delta P_{i}=\gamma_{0}^{\rm bose}\gamma_{i}^{\rm bose}$. This gives, in the same small $\mathbf{k}$ limit,
@@
with eigenvalues 00 and ±$\pm ik$, respectively, where $k=|\mathbf{k}|=\sqrt{k_{X}^{2}+k_{Y}^{2}+k_{Z}^{2}}$. These correspond to a longitudinal mode, and right- and left-polarized modes, respectively. The Hamiltonian thus has two zero-energy longitudinal eigenstates, as well as a positive- and negative-energy eigenstate for each circular polarization. In the free theory, these do not mix, so if we consider positive-energy initial states, the dynamics do not take us out of the positive-energy subspace. That is not generally true in the presence of interactions, which we will discuss later.
When we promote the single-particle theory to allow for the creation and annihilation of photons, we will find that the field operators, expressed in the usual manner in terms of the creation and annihilation operators and the above eigenvectors, will satisfy the Maxwell equations in the long-wavelength limit. To see that here, for the one-particle sector, we have to consider the photon wave function (see Refs. BialynickiBirula ; Darwin32 ):
@@
and the vector identity in Equation (25), we obtain
@@
Substituting $-i\boldsymbol{\nabla}$ for $c\mathbf{k}$ and considering the real and imaginary parts separately, we obtain the two time-dependent Maxwell equations:
@@
The other two Maxwell equations (in the absence of charge and current) arise from the transversality of the positive-energy eigenvectors.
@@ 5.1 Fermions
As was shown in MlodinowBrun21 , it is possible to promote the single-particle fermionic theory to a multi-particle QCA that reduces to the usual Dirac field theory in the long-wavelength limit. To accomplish this, one begins with a maximum number of particles, $N_{\rm max}$, and then lets $N_{\rm max}$ become arbitrarily large. In other words, one begins with the space
@@
The subspace $\mathcal{A}_{n}$ is the totally antisymmetric subspace with $n$ particles, and $\mathcal{H}_{0}$ is the one-dimensional vacuum space ${\rm span}({|\Omega\rangle})$, where ${|\Omega\rangle}={{|\omega\rangle}}^{\otimes N_{\rm max}}$. We can define basis vectors for the subspaces $\mathcal{A}_{n}$ in terms of the “energy eigenstates” of the quantum walk:
@@
The $n$-particle antisymmetrized subspace is thus preserved by the time evolution, and, hence, so is the full physical subspace $\mathcal{H}_{\rm phys}$: an antisymmetric initial state will evolve to be antisymmetric at all later times. Given this basis, we can define a set of creation and annihilation operators that transform the energy basis vectors into each other:
@@
and that obey the usual anti-commutation relations. To avoid redundant counting of the basis vectors, we should establish a standard ordering for the states $\mathbf{k}_{i},j_{i}$. The choice of this ordering is unimportant, so long as it is unambiguous.
@@
A very important observation should be made here. These creation and annihilation operators $a^{\dagger}_{\mathbf{k},j}$ and $a_{\mathbf{k},j}$ are clearly nonlocal operators; so the unitary operator on the right-hand side of Equation (37) is nonlocal. But the unitary operator $U_{\rm total}$ is itself purely local. It is a collection of local quantum walk unitaries and can easily be embedded in a quantum cellular automaton. These two sides are equal only on states in the physical subspace. Indeed, the right-hand side has not been defined for states outside that subspace.
@@
using the matrices $V^{\dagger}_{\mathbf{k}}$ from Equation (16). These vectors have an effective time evolution
@@
where $U_{\mathbf{k}}$ is given by Equation (16).
What about the maximum number of particles $N_{\rm max}$? This may seem like an unphysical feature of the theory. In the free theory, particle number is conserved, so a solution will generally be confined to a single antisymmetric subspace $\mathcal{A}_{n}$. However, in the interacting case, this conservation need not hold.
One way to handle this is to start with a vacuum corresponding to a Dirac Sea, in which all the negative-energy states are occupied, while all positive-energy states are vacant. Interactions do not create or destroy particles but may move a particle from a negative- to a positive-energy state, leaving a hole (corresponding to particle–anti-particle pair creation), or move a positive-energy state to a negative-energy hole (corresponding to particle–anti-particle annihilation). In such a construction, we would have $N_{\rm max}=2N^{3}$, where $N$ is the lattice size. This means that the Hilbert spaces of the local subsystems at each lattice site are enormously high-dimensional (albeit finite), but the states of the theory remain restricted to the much lower-dimensional physical subspace. In the full space, the dynamics of the free theory are fully local; the effective dynamics in the physical subspace obey Fermi statistics.
@@ 5.2 Bosons
We can make a similar construction for free bosons by taking $N_{\rm max}$ copies of a bosonic QW, representing distinguishable particles, and then confining ourselves to the completely symmetric subspace. So long as the total number of bosonic particles $N$ is less than $N_{\rm max}$, this system will behave like a bosonic field theory with $N$ excitations, and we can introduce creation and annihilation operators and rewrite the evolution unitary in terms of them, just as we did in the fermionic case.
@@
where $j=X,Y,Z$. As discussed in Section 4, we can recover a Maxwell field if we use a doubled space, where each of the projectors is two-dimensional, for an internal space dimension of six. To simplify the derivation of the corresponding QCA, we will first consider a non-doubled space, with an internal dimension three, so we can write the projectors as outer products of the eigenstates of the $J_{j}$ matrix:
@@
In other words, the shifts are performed in two steps: the $+$ and -- states at neighboring sites are interchanged, and then at each site, $+$ and -- are switched back. (The 0 states remain unchanged throughout.)
The advantage of breaking this down into two steps is that each step is a purely local operation that only affects two neighboring sites. This makes it very easy to turn into a QCA that is definitely local. In constructing the QCA, each basis state ${|\mathbf{x},\varepsilon\rangle}$ of the QW corresponds to a harmonic oscillator mode with creation and annihilation operators $a^{\dagger}_{\mathbf{x},\varepsilon},a_{\mathbf{x},\varepsilon}$. Then, the evolution unitary becomes
@@
We can see that $\hat{U}_{j}$ is a product of tensor products of unitaries, which act on the local subsystems $\mathbf{x}$ or on non-overlapping sets of modes at neighboring subsets $\mathbf{x}$ and $\mathbf{x}+\Delta x\mathbf{\hat{j}}$. So $\hat{U}_{j}$ is local and defines a QCA. Since operators on different modes commute, we can rewrite each tensor product of exponentials as the exponential of a sum of terms (where now we understand the creation and annihilation operators as acting as the identity on other modes). Further, we see that the defined QCA preserves the particle number (or the number of excitations); and on the single-particle sector, it exactly reproduces a quantum walk.
@@
Each of the operators $\hat{U}_{X,Y,Z}$ has the same form as in Equation (42). In the quantum walk, we obtain nontrivial evolution because the projectors $P^{\pm}_{X,Y,Z}$ for the shifts along each axis project onto subspaces that are non-orthogonal to those of the other axes. For the QCA, this means that the creation and annihilation operators in Equation (42) for one axis $j$ are linear combinations of those for a different axis $j^{\prime}$.
Since all the expressions inside the exponentials are quadratic in creation and annihilation operators, we can rewrite Equation (42) in terms of a standard set of modes with operators $a^{\dagger}_{\mathbf{x},\ell}$ and $a_{\mathbf{x},\ell}$:
@@
It is easy to check that creation and annihilation operators for different values of $\mathbf{k}$ commute with each other. If we substitute this definition into Equation (44), the unitary decomposes into a product of factors acting on different momentum modes:
@@
This is most readily achieved by diagonalizing $A$ and $B$, exponentiating them, and then multiplying them. One can find $C$ by diagonalizing the resulting matrix and taking its log. In the case of the unitary factor $U_{j}$, the matrices $A_{j}$ and $B_{j}$ have elements
@@
These matrices are $3\times 3$, so, in general, it is not difficult to diagonalize them. For a bosonic QW, we suggested in Section 3 that the basis vectors should be eigenvectors of the three spin-1 $J$ matrices given in Equation (12). This choice for the bosonic QCA would lead to three pairs of $A$ and $B$ matrices:
@@
where we have used the compact notation $c_{x,y,z}=\cos(k_{x,y,z}\Delta x)$ and $s_{x,y,z}=\sin(k_{x,y,z}\Delta x)$.
This matrix can, in principle, be diagonalized in closed form, but the expression is dauntingly complicated. Its eigenvalues can also be calculated, though they are also somewhat complicated:
@@
Simple observation shows that the $+$ and -- eigenvalues have opposite phases: $\phi_{-}=-\phi_{+}$. These can be thought of as negative- and positive-energy modes. We can recover more familiar expressions by going to the long-wavelength limit, where $|\mathbf{k}\Delta x|\ll 1$. In this limit,
@@
As discussed in Section 4, we can recover the Maxwell field theory by doubling the internal space. In place of the spin matrices $J_{j}$ we use the $6\times 6$ matrices $\sigma_{Z}\otimes J_{j}$. The derivation above is essentially doubled, giving two zero-energy, two negative-energy, and two positive-energy modes for each allowed momentum $\mathbf{k}$. If we confine ourselves to the positive-energy subspace, these two modes correspond to the two allowed helicity states of the electromagnetic field. The two negative-energy modes are like retarded versions of these two modes. The zero-energy modes are static and unaffected by the other modes.
We have seen that the free theories of both fermionic and bosonic QCAs include matching positive- and negative-energy modes (and zero-energy modes as well, in the bosonic case). This turns out to be a general property of QCAs based on their overall symmetry, as discussed in the next section.